A car has a mass of 1500 kg moves with a speed of 40 m/s. If the car speeds up to 50 m/s. How much is the work done by the car engine?


1
Expert's answer
2022-01-27T09:11:37-0500

Given:

"m=1500\:\rm kg"

"v_1=40\:\rm m/s"

"v_2=50\:\rm m/s"


The energy-work theorem says

"W=\Delta E=\frac{mv_2^2}{2}-\frac{mv_1^2}{2}"

"W=\frac{1500*(50)^2}{2}-\frac{1500*(40)^2}{2}=675 000\: \rm J"